# aims-PAX: Parallel Active eXploration for the automated construction of Machine Learning Force Fields

### v1

@@ Abstract
Recent advances in machine learning force fields (MLFF) have significantly extended the capabilities of atomistic simulations. This progress highlights the critical need for reliable reference datasets, accurate MLFFs, and, crucially, efficient active learning strategies to enable robust modeling of complex chemical and materials systems. Here, we introduce aims-PAX, an automated, multi-trajectory active learning framework that streamlines the development of MLFFs. Designed for both experts and newcomers, aims-PAX offers a modular, high-performance workflow that couples flexible sampling with scalable training across CPU and GPU architectures. Built on the widely adopted ab initio code FHI-aims, the framework seamlessly integrates with state-of-the-art ML models and supports pretraining using general-purpose (or "foundational") models for rapid deployment in diverse systems. We demonstrate the capabilities of aims-PAX on two challenging systems: a highly flexible peptide and bulk CsPbI3 perovskite. Across these cases, aims-PAX achieves a reduction of up to two orders of magnitude in the number of required reference calculations and enables over 20x speedup in AL cycle time through optimized resource utilization. This positions aims-PAX as a powerful and versatile platform for next-generation ML-driven atomistic simulations in both academic and industrial settings.
@@ I Introduction:
The success of Machine Learning Force Fields (MLFFs)Unke et al. (2021) has deeply transformed the field of molecular simulations. They are now the preferred method for simulating the dynamics of large systems, such as extensive perovskite unit cellsBaldwin et al. (2024) or solvated proteinsUnke et al. (2024), with quantum-chemical accuracy. While general-purpose (GP) (sometimes called “foundational”) modelsKabylda et al. (2025); Smith, Isayev, and Roitberg (2017); Kovács et al. (2025); Chen et al. (2019); Deng et al. (2023a); Choudhary and DeCost (2021); Merchant et al. (2023); Gasteiger, Becker, and Günnemann (2021) trained on large datasetsHoja et al. (2021); Eastman et al. (2023); Deng et al. (2023b); Barroso-Luque et al. (2024); Levine et al. (2025); Ganscha et al. (2025) are becoming more widespread, there remains a strong demand for high-quality data to fine-tune these models or to build new custom models for challenging applications.Novelli et al. (2024); Kaur et al. (2025); Radova et al. (2025)
@@
The process of collecting representative high-quality datasets can be labor-intensive, requiring considerable manual effort and computational resources. To address these challenges, a common approach is to employ active learning (AL).Settles (2009); Ren et al. (2021) In AL, an uncertainty measure of a model’s prediction is used to select data points for labeling and inclusion in the training dataset. This approach enriches the training dataset with data points that represent a challenge for the current state of the model. In essence, the model autonomously determines which data to prioritize for training and which to disregard. Therefore, this procedure reduces human intervention and decreases the computational cost of model training by requiring only a small number of costly and slow reference ab initio calculations to reach an acceptable accuracy. In addition, it also improves the robustness of the MLFF by auto-detecting and correcting its possible failures during the training procedure.
@@
AL has been successfully applied to a plethora of applications. For example, Young et al.Young et al. (2021) used active learning to iteratively improve a MLFF that was able to accurately simulate solvents and selected chemical reactions. In a study by Stark et al.Stark et al. (2023), an AL workflow leveraging clustering algorithms was used to model reactive hydrogen dynamics on copper surfaces. Furthermore, Mohanty et al.Mohanty et al. (2023) showed how AL was necessary to augment a dataset for efficiently training MLFFs for polymer dynamics and Kang et al.Kang et al. (2024a) highlighted how AL was crucial to model strongly anharmonic materials. Numerous other successful AL applications can be found in the literature. Zhang et al. (2019a); Smith et al. (2018); Zhang et al. (2023); Shambhawi, Csányi, and Lapkin (2021); Sivaraman et al. (2020); Kuryla et al. (2025); Erhard et al. (2024); Vandermause et al. (2024); Duschatko et al. (2024); Xie et al. (2023); Vandermause et al. (2022); Johansson et al. (2022); Xie et al. (2021); Matsumura et al. (2025); Gurlek et al. (2025) While AL is always beneficial in the data collection process, the automation degree of the procedure varies broadly. Often, AL is done manually or by users who develop tailored scripts for their specific problems. This situation results in the need for expert knowledge, such as selecting starting geometries, setting uncertainty thresholds, or deciding when to stop sampling. Additionally, employing collections of custom scripts instead of a defined workflow makes the process less accessible to new practitioners and less reproducible by other researchers. In recent years, the community has started addressing these challenges by offering various automated software solutions. For example, in the DFT codes such as the Vienna ab initio simulation package (VASP)Kresse and Hafner (1993); Kresse and Furthmüller (1996a, b), CASTEPClark et al. (2005); Stenczel et al. (2023) and the Amsterdam Modeling Suite (AMS)Rüger et al. (2025) different automated AL workflows are implemented. Next to AL methods directly integrated into quantum chemistry codes, there also exist separate software packages offering AL or automated simulation functionalities such as FLAREVandermause et al. (2020), CatFlowLiu et al. (2025), AlebrewZaverkin et al. (2024), PsiFlowVandenhaute et al. (2023), ALmoMDKang et al. (2024b), apaxSchäfer et al. or PALZhou et al. (2024). Although such tools have helped establish MLFFs and AL as a standard tool in molecular simulations, there is a potential for improvements that we address in this work, in particular with respect to the efficiency of configurational space exploration, hardware utilization and support for multisystem sampling.
@@
We present aims-PAX, short for ab initio molecular simulation-Parallel Active eXploration, a flexible, fully automated, open-source software package for performing AL using a parallelized algorithm that enables efficient resource management. The current implementation is coupled to the FHI-aimsAbbott et al. (2025) program for DFT calculations and the MACEBatatia et al. (2022, 2025) architecture as an MLFF model. However, the algorithm itself is agnostic to both the MLFF architecture and the DFT code.
We showcase aims-PAX performance by applying it to a large, flexible peptide. In the case of this peptide, aims-PAX automatically produces accurate and stable MLFFs and provides a high-quality dataset with a minimal number of ab initio calculations. For the latter, we compare aims-PAX performance to a more traditional manual data collection strategy used in the TEA challenge.Poltavsky et al. (2025a, b) Fully automated, aims-PAX reduced the number of required DFT calculations by two orders of magnitude while achieving essentially the same results.
The structure of the article is the following. First, in the Methods section we provide a technical description of the aims-PAX workflow, as well as a thorough report of computational methods and their settings used. In the Results section, we report the application of aims-PAX to the peptide Ac-F-A5-K. In addition, the computational advantages of the parallelized AL algorithm are shown on the example of bulk CsPbI3 pervoskite. We conclude the article with a short summary and outlook for the future of aims-PAX.
II Methods
In this section, we describe the proposed aims-PAX AL workflow by first explaining how initial datasets and initial models can be generated efficiently through an integrated procedure that leverages GP-MLFFs. Afterwards, we explain the core of the proposed method, which consists of the parallelized AL algorithm. We describe the implementation details, multi-trajectory sampling, and the employed uncertainty measure, along with its adaptive threshold. We close the section with a comprehensive account of the computational methods used in the various studies.
@@ II.1 Initial Dataset and Model Generation
The described initial dataset generation (IDG) can be established using one of two approaches:
@@
The second approach is generally preferable, as it helps decorrelate the geometries, making the IDG significantly more computationally efficient—by at least an order of magnitude. Also, we coupled the IDG with ParslBabuji et al. (2019), similar to what is done in PsiflowVandenhaute et al. (2023). This enables users to perform DFT calculations on sampled geometries across multiple nodes in parallel. Importantly, the GP model does not need to provide accurate energies or forces; it acts solely as a geometry generator in combination with MD simulations.
Both initialization strategies are implemented in aims-PAX, see Fig. 1b. Currently, the implementation includes the MACE-MP0Batatia et al. (2024) model, with additional models to be incorporated in the future, such as SO3LR Kabylda et al. (2025).
Once the dataset reaches a user-defined threshold in size, it is split into several equally sized subsets. These subsets are randomly selected from the full dataset, with one subset being assigned to each MLFF ensemble member. This ensures that each model is trained and validated on slightly different data. In addition to varying the training data, we introduce further diversity between ensemble members by using different random seeds for initializing model weights. Each MLFF is then trained on its assigned subset for a user-specified number of epochs.
Optionally, this procedure of generating data and training can be repeated multiple times. More precisely, after training, new structures are sampled and subsequently labeled, which is followed by more training steps. More details on this approach can be found in Section A.3.1.
During these cycles, the MLFFs are trained without reinitializing their weights. Instead, the existing weights are reused at every training step, and models are trained on their entire datasets to prevent catastrophic forgettingFrench (1999); McCloskey and Cohen (1989); McClelland, McNaughton, and O’Reilly (1995); Ratcliff (1990); Kumaran, Hassabis, and McClelland (2016). This continual learning (CL) approach Kirkpatrick et al. (2017); Aljundi (2019); Parisi et al. (2019) enables models to improve iteratively without retraining from scratch each time, reducing the number of required training steps. Crucially, at this stage, models are not trained to full convergence; instead, training is deliberately limited to a small number of epochs (typically five or fewer). The described early-stopping strategy avoids overfitting the models on the initial datasets and hinders them from getting stuck in local minima. This would make updating the models with new data during the AL significantly more difficult without reinitializing their weights.
@@ II.2 Parallelized Active Learning
In the case of aims-PAX, each time the threshold is crossed and the DFT calculation has been performed, the new data is added to the training (or validation) set of all MLFFs. These are then updated in a CL scheme using a user-specified, ideally low, number of epochs similarly to the one employed in the IDG. For more details on the exact training strategy we refer to Section A.3.1.
While the algorithm proposed herein is, in principle, agnostic to the choice of uncertainty quantification method, we employ the query by committee (QBC)Seung, Opper, and Sompolinsky (1992); Schran, Brezina, and Marsalek (2020); Artrith and Behler (2012) approach due to its conceptual simplicity and widespread adoption. The integration of alternative uncertainty estimation techniques into our framework is straightforward and will be explored in future work.
In the QBC approach, an ensemble of independently trained ML models is used to produce a distribution of predicted outputs during inference. As described previously, diversity among ensemble members arises from differences in initial weight initialization seeds and distinct initial training datasets. The variance within the ensemble predictions serves as a way to quantify a model’s uncertainty. Specifically, we quantify uncertainty based on the variance of atomic force predictions, using the maximum per-atom force variance across the system, as defined in Eq. 1Zhang et al. (2019a),
@@
The primary advantage of this adaptive thresholding approach is that it eliminates the need for a fixed, user-defined uncertainty cutoff, which can vary between systems.Kulichenko et al. (2023) Since the moving average naturally decreases over time, some configurations will always exceed the threshold. As a result, the sampling frequency depends on the value of $c_{x}$: if set too high, very few points may be sampled; if too low, the method may oversample. Based on our experience and also reported for the MLFF training in the VASP code, values of $c_{x}\in\left[-0.1,\,0.1\right]$ serve as practical starting points.
To improve the efficiency and robustness of the active sampling, we adopt a multi-trajectory approach that has also been successfully applied in similar frameworks.Liu et al. (2025); Zhou et al. (2024); Matsumura et al. (2025) Herein multiple ML-driven simulations are executed in parallel, see Fig. 1c. These trajectories may differ in their sampling strategies, utilizing various thermostats, barostats, external conditions, or simulation schemes ranging from classical MD to enhanced sampling techniques such as metadynamics. Although only a subset of these options is currently implemented, additional methods are under development and will be integrated in future updates. We point out that the uncertainty threshold as defined in Eq. 2 is shared across all of these trajectories.
@@
Similarly to the IDG, ParslBabuji et al. (2019) can also be used here, allowing to distribute DFT calculations across multiple nodes in parallel. Additionally, the number of DFT workers can be adapted dynamically up to a user-defined maximum. Through the flexible allocation of resources, we ensure that no worker is idle or overloaded. This is particularly useful in AL as the demand for new data can change during the procedure. For example, there can be long sequences where an MLFF trajectory is certain about all encountered geometries or intervals where all trajectories lead to high uncertainty regions.

### v2

# aims-PAX: Parallel Active eXploration Enables Expedited Construction of Machine Learning Force Fields for Molecules and Materials

@@ Abstract
Recent advances in machine learning force fields (MLFF) have significantly extended the reach of atomistic simulations. Continuous progress in this field requires reliable reference datasets, accurate MLFF architectures, and efficient active learning strategies to enable robust modeling of complex molecular and material systems. Here we introduce aims-PAX, an expedited, multi-trajectory active learning framework that streamlines the development of stable and accurate MLFFs. Designed for a wide range of researchers, aims-PAX offers a modular, high-performance workflow that couples diversified sampling with scalable training across CPU and GPU architectures. Integrated with the widely used ab initio code FHI-aims, the framework supports state-of-the-art ML models and dataset generation using general-purpose (or "foundational") force-fields for rapid deployment in diverse systems. We demonstrate the capabilities of aims-PAX in various challenging tasks: creating datasets and models for highly flexible peptides, multiple organic molecules at once, explicitly solvated molecules, and for efficiently handling computationally demanding systems such as the CsPbI3 perovskite. We show that aims-PAX achieves a reduction of up to three orders of magnitude in the number of required reference calculations, automatically selects challenging systems within a given chemical space, facilitates simulation of solvated molecules with more than thousand atoms, while enabling a ten-fold speedup in active-learning time through optimized resource utilization. This positions aims-PAX as a powerful and versatile platform for next-generation atomistic simulations in both academic and industrial settings.
@@ I Main:
The successes of machine learning force fields (MLFFs)Unke et al. (2021) have deeply transformed the field of molecular simulations. They are now the preferred method for simulating the dynamics of large systems, such as perovskites Baldwin et al. (2024) or solvated proteins Unke et al. (2024), with quantum-chemical accuracy. While general-purpose (GP) (sometimes called “foundational”) models Kabylda et al. (2025); Smith, Isayev, and Roitberg (2017); Kovács et al. (2025); Chen et al. (2019); Deng et al. (2023a); Choudhary and DeCost (2021); Merchant et al. (2023); Gasteiger, Becker, and Günnemann (2021) trained on large datasets Hoja et al. (2021); Eastman et al. (2023); Deng et al. (2023b); Barroso-Luque et al. (2024); Levine et al. (2025); Ganscha et al. (2025) are becoming more widespread, there remains a strong demand for high-quality data to fine-tune these models or to build new, and often cheaper, custom models for challenging applications. Novelli et al. (2024); Kaur et al. (2025); Radova et al. (2025)
@@
The process of collecting representative high-quality datasets can be labor-intensive, requiring considerable manual effort and computational resources. To address these challenges, a common approach is to employ active learning (AL). Settles (2009); Ren et al. (2021) In AL, an uncertainty measure of a model prediction is used to select data points for labeling and inclusion in the training dataset. This approach enriches the training dataset with points that represent a challenge for the current state of the model. In essence, the model autonomously determines which data to prioritize for training and which to disregard. Therefore, this procedure reduces human intervention and decreases the computational cost of model training by requiring only a small number of expensive and slow reference ab initio calculations to reach an acceptable accuracy. In addition, AL also improves the robustness of the MLFF by detecting and correcting possible failures during the training procedure.
AL has been successfully applied to a plethora of applications. For example, Young et al. Young et al. (2021) used active learning to iteratively improve a MLFF that was able to accurately simulate solvents and selected chemical reactions. In a study by Stark et al. Stark et al. (2023), an AL workflow leveraging clustering algorithms was used to model reactive hydrogen dynamics on copper surfaces. Furthermore, Mohanty et al. Mohanty et al. (2023) showed how AL was necessary to augment a dataset for efficiently training MLFFs for polymer dynamics and Kang et al. Kang et al. (2024a) highlighted how AL was crucial to model strongly anharmonic materials. Numerous other successful AL applications can be found in the literature. Zhang et al. (2019); Smith et al. (2018); Zhang et al. (2023); Shambhawi, Csányi, and Lapkin (2021); Sivaraman et al. (2020); Kuryla et al. (2025); Erhard et al. (2024); Vandermause et al. (2024); Duschatko et al. (2024); Xie et al. (2023); Vandermause et al. (2022); Johansson et al. (2022); Xie et al. (2021); Matsumura et al. (2025); Gurlek et al. (2025)
@@
While AL is always beneficial in the data collection process, the automation degree of the procedure varies broadly. Often, AL is done manually or by users who develop tailored scripts for their specific problems. This situation results in the need for expert knowledge, such as selecting starting geometries, setting uncertainty thresholds, or deciding when to stop sampling. Additionally, employing collections of custom scripts instead of a defined workflow makes the process less accessible to new practitioners and less reproducible by other researchers. In recent years, the community has started addressing these challenges by offering various automated software solutions. For example, in the DFT codes such as the Vienna ab initio simulation package (VASP)Kresse and Hafner (1993); Kresse and Furthmüller (1996a, b), CASTEPClark et al. (2005); Stenczel et al. (2023) and the Amsterdam Modeling Suite (AMS)Rüger et al. (2025) different automated AL workflows are implemented. Next to AL methods directly integrated into quantum chemistry codes, there also exist separate software packages offering AL or automated simulation functionalities such as FLAREVandermause et al. (2020), CatFlowLiu et al. (2025), AlebrewZaverkin et al. (2024), PsiFlowVandenhaute et al. (2023), ALmoMDKang et al. (2024b), apaxSchäfer et al. or PALZhou et al. (2024). Although such tools have helped establish MLFFs and AL as a standard tool in molecular simulations, there is a potential for improvements that we address in this work, in particular with respect to the efficiency of configurational space exploration, hardware utilization, support for multi-system sampling and seamless data generation for periodic materials and finite molecular systems.
@@
We present aims-PAX, short for ab initio molecular simulation-Parallel Active eXploration, as a fully automated open-source software package for performing AL, using a parallelized algorithm that enables efficient resource management. The current implementation is integrated with the FHI-aims Abbott et al. (2025) program for DFT calculations and the MACE Batatia et al. (2022, 2025) architecture as an MLFF model. However, while the software is developed primarily for working in tandem with the codes and models named above, the algorithm itself is agnostic to both the MLFF architecture and the choice of DFT code.
Importantly, we want to highlight defining features of aims-PAX that emphasize its versatility and uniqueness:
  1. 1.
    Leverage of GP-MLFFs for data acquisition
  2. 2.
    Multi-system sampling for transferable MLFFs
  3. 3.
    Flexible combination of ab initio levels of theory
  4. 4.
    Seamless handling of molecular and materials systems
  5. 5.
    Efficient CPU/GPU workload management
  6. 6.
    Support for state-of-the-art neural network based MLFFs
We demonstrate the capabilities of aims-PAX herein on a flexible peptide, where it autonomously generates accurate and stable MLFFs using two orders of magnitude fewer DFT calculations than traditional workflows.Poltavsky et al. (2025a, b) Beyond individual systems, aims-PAX concurrently samples multiple molecules, autonomously identifying the most informative configurations to build a single, transferable MLFF that generalizes across chemical space. Integrated seamlessly with FHI-aims, it unifies gas-phase, bulk, and solvated regimes—capturing paracetamol in vacuum, microsolvated by water, and explicit solvent within one model. Finally, large-scale tests on a bulk perovskite demonstrate its exceptional scalability and computational efficiency, establishing aims-PAX as a universal framework for automated, data-driven force-field generation.
II Results
Figure 1: Overview of the aims-PAX workflow: (a) Required input files: The first file (control.in) follows FHI-aims conventionsBlum et al. (2009) and contains the DFT settings. It is also possible to use different DFT settings per trajectory. The system’s geometry, or initial geometries can either be inside a folder (geometry/) or, in the case of a single geometry, in a file (geometry.in). The file (mace.yaml) contains MACEBatatia et al. (2022, 2025) model hyperparameters and the fourth (aims_PAX.yaml) is an aims-PAX-specific file containing the IDG and AL settings. For the AL workflow, folders containing the initial datasets (data/) and models (models/) are required. (b) Initial dataset generation (IDG): Geometries are sampled using either DFT or a GP model, with DFT providing labels in both cases. Sampling continues until a (* user specified) criterion is met. (c) Parallelized active learning: The AL workflow requires input files, existing data, and models, which can be provided by the IDG procedure. Sampling occurs over multiple trajectories, triggering DFT calculations when an uncertainty threshold is exceeded. GPU-based ML tasks (orange) and CPU-based DFT tasks (blue) can run in parallel. AL is continued until a (* user specified) stopping condition is met. (d) Output: Models and collected data produced during AL (and IDG).
@@ II.1 Initial Dataset and Model Generation
Such an initial dataset generation (IDG) can, for example, be established using one of two approaches:
@@
The second approach is generally preferable, as it helps decorrelate the geometries, making the IDG significantly more computationally efficient. Both initialization strategies are implemented in aims-PAX, see Fig. 1b. We also couple the IDG with ParslBabuji et al. (2019), similar to what is done in PsiflowVandenhaute et al. (2023). This enables users to perform DFT calculations on sampled geometries across multiple nodes in parallel. Importantly, the GP model does not need to provide accurate energies or forces; it acts solely as a geometry generator in combination with MD simulations. Currently, the implementation includes the MACE-MP Batatia et al. (2024) and SO3LR Kabylda et al. (2025) GP models, with additional models to be incorporated in the future.
Once the dataset reaches a user-defined threshold in size, it is split into several equally sized subsets. These subsets are randomly selected from the full dataset, with one subset being assigned to each MLFF ensemble member. This ensures that each model is trained and validated on slightly different data. In addition to varying the training data, we introduce further diversity between ensemble members by using different random seeds for initializing model weights. Each MLFF is then trained on its assigned subset for a user-specified number of epochs. Optionally, this procedure of generating data and training can be repeated multiple times. More precisely, after training, new structures are sampled and subsequently labeled, which is followed by more training steps. More details on this approach can be found in Section S5A.
During these cycles, the MLFFs are trained without reinitializing their weights. Instead, the existing weights are reused at every training step, and models are trained on their entire datasets to prevent catastrophic forgettingFrench (1999); McCloskey and Cohen (1989); McClelland, McNaughton, and O’Reilly (1995); Ratcliff (1990); Kumaran, Hassabis, and McClelland (2016). This continual learning (CL) approach Kirkpatrick et al. (2017); Aljundi (2019); Parisi et al. (2019) enables models to improve iteratively without retraining from scratch each time, reducing the number of required training steps. Crucially, at this stage, models are not trained to full convergence; instead, training is deliberately limited to a small number of epochs. The described early-stopping strategy avoids overfitting the models on the initial datasets and hinders them from getting stuck in local minima. This would make updating the models with new data during the AL significantly more difficult without reinitializing their weights.
@@ II.2 Parallelized Active Learning
In the case of aims-PAX, each time the threshold is crossed and the DFT calculation has been performed, the new data is added to the training (or validation) set of all MLFFs. These are then updated in a CL scheme using a user-specified, ideally low, number of epochs similarly to the one employed in the IDG. For more details on the exact training strategy we refer to Section S5A. While the algorithm proposed herein is, in principle, agnostic to the choice of uncertainty quantification method, we employ the query by committee (QBC)Seung, Opper, and Sompolinsky (1992); Schran, Brezina, and Marsalek (2020); Artrith and Behler (2012) approach due to its conceptual simplicity and widespread adoption. The integration of alternative uncertainty estimation techniques into our framework is straightforward and will be explored in future work.
In the QBC approach, an ensemble of independently trained ML models is used to produce a distribution of predicted outputs during inference. As described previously, diversity among ensemble members arises from differences in initial weight initialization seeds and distinct initial training datasets. The variance within the ensemble predictions serves as a way to quantify a model’s uncertainty. Specifically, we quantify uncertainty based on the variance of atomic force predictions, using the maximum per-atom force variance across the system, as defined in Eq. 1Zhang et al. (2019),
@@
The primary advantage of this adaptive-threshold approach is that it eliminates the need for a fixed, user-defined uncertainty cutoff, which can vary between systems.Kulichenko et al. (2023) Since the moving average naturally decreases over time, some configurations will always exceed the threshold. As a result, the sampling frequency depends on the value of $c_{x}$: if set too high, very few points may be sampled; if too low, the method may oversample. Based on our experience and also reported for the MLFF training in the VASP code, values of $c_{x}\in\left[-0.1,\,0.1\right]$ serve as practical starting points.
To improve the efficiency and robustness of the active sampling, we adopt a multi-trajectory approach that has also been successfully applied in similar frameworks.Liu et al. (2025); Zhou et al. (2024); Matsumura et al. (2025) Herein multiple ML-driven simulations are executed in parallel, see Fig. 1c. These trajectories may differ in their sampling strategies, utilizing various thermostats, barostats, external conditions, or simulation schemes. Importantly, different trajectories can also simulate different systems. We point out that the uncertainty threshold as defined in Eq. 2 is shared across all of these trajectories.
@@
Similarly to the IDG, ParslBabuji et al. (2019) can also be used here, allowing to distribute DFT calculations across multiple nodes in parallel. Additionally, the number of DFT workers can be adapted dynamically up to a user-defined maximum. Through the flexible allocation of resources, we ensure that no worker is idle or overloaded. This is particularly useful in AL as the demand for new data can change during the procedure. For example, there can be long sequences where an MLFF trajectory is certain about all encountered geometries or intervals where all trajectories lead to high uncertainty regions. Due to these crucial computational benefits this approach has been designated as the default procedure in aims-PAX.